# Data-Centric Machine Learning for Geospatial Remote Sensing Data

### v3

@@ Abstract
keywords:
data-centric machine learning , data quality ,  review ,  data curation ,  data valuation
[…]
[inst3]organization=Laboratory of Geo-information Science and Remote Sensing, Wageningen University,addressline=Droevendaalsesteeg 3, city=Wageningen, postcode=6708 PB, state=Gelderland, country=Netherlands
[…]
[inst4]organization=Department of Earth Observation Science, Faculty ITC, University of Twente,addressline=Drienerlolaan 5, city=Enschede, postcode=7522 NB, state=Overijssel, country=the Netherlands
[…]
[inst5]organization=Department of Physics and Technology, UiT The Arctic University of Norway,addressline=Klokkargårdsbakken 35, city=Tromsø, postcode=9019, country=Norway
[…]
[inst6]organization=Department of Computer Science, University of Sheffield,addressline=211 Portobello, Sheffield City Centre, city=Sheffield, postcode=S1 4DP, country=United Kingdom
[…]
[inst7]organization=MICS Laboratory, CentraleSupélec, Paris-Saclay University,addressline=3 Rue Joliot Curie, city=Gif-sur-Yvette, postcode=91190, country=France
[…]
[inst8]organization=Microwaves and Radar Institute, German Aerospace Center (DLR),addressline=Münchner Str. 20, city=Oberpfaffenhofen-Wessling, postcode=82234, country=Germany
[…]
[inst7]organization=Computing Science and Mathematics, University of Stirling,city=Stirling, postcode=FK9 4LA, country=United Kingdom
[…]
[inst9]organization=Department of Aerospace Engineering, University of the Bundeswehr Munich,addressline=Werner-Heisenberg-Weg 39, city=Neubiberg, postcode=85579, state=Bavaria, country=Germany
[…]
[inst10]organization=Environmental Computational Science and Earth Observation Laboratory, Ecole Polytechnique Fédérale de Lausanne (EPFL),addressline=Ronquos 86, city=Sion, postcode=1951, country=Switzerland
1 Introduction
@@
Figure 1: Steps of the machine learning cycle. Each step highlights one way to interact with data and its quality, and each step can employ multiple techniques to perform the interaction (see Section 3). It involves problem definition, data creation and curation, model training and evaluation, and final deployment, which feeds back into a modified problem definition. Model-centric learning focuses primarily on model training and evaluation, while data-centric learning involves algorithms covering data curation, creation, specific training strategies, evaluation, and deployment feedback. The considered quality criteria are diversity and completeness, accuracy, consistency, unbiasedness, and relevance, see Section 2).
@@
The article is organized as follows: It first clarifies the terminology, definitions, and goals of data-centric and model-centric learning (Section 2) and then provides a comprehensive discussion of data-centric machine learning and its steps, with a focus on the analysis and interpretation of geospatial data. Additionally, it presents a systematic review of data-centric machine learning techniques tailored explicitly to geospatial data (Section 3), facilitating the learning and evaluation of high-performing machine learning models (see Fig. 1). The paper is supported by three experiments using data-centric machine learning techniques, presented in Section 4.
2 Definitions and Terminology
@@
2.1 Machine learning cycle
@@
2.2 High-quality criteria for geospatial data
@@
3 Data-centric machine learning techniques
@@
Step 5: evaluationuse quality informationStep 4: model traininguse quality informationStep 3: curationact on qualityStep 2: creationact on qualitydiversity & completenessaccuracyconsistencyunbiasednessrelevance Train/eval splits [40, 41] Resampling techniques [42] Slicing [43] Out-of-distribution evaluation [44, 45] Adversarial attacks [46] Sensitivity analysis [47, 48] Curriculum learning [49, 50, 51, 52, 53, 54] Quality-based weighted loss [55, 56] Pre-training [57, 58, 59, 60, 61] Continual learning [62] Missing data imputation and inpainting [63, 64, 65, 66, 67, 68] Core-set selection [69, 70] Label noise reduction and confident learning[71] Multi-labeler error detection [72] Bias mitigation [73] Dimensionality reduction [74, 75, 76] Self-training [77] Augmentation [78, 79, 80, 81] Collection and sampling strategies [82, 83, 84, 58, 85, 86, 87, 88, 85, 89, 90] Active learning [91, 92, 93] Confident learning [94] Multi-labeler agreement assessment [33, 95] Bias detection and mitigation [27] Relevance and influence assessment [96, 90]
@@
3.1 Data Creation (Step 2)
@@
To quantify the relevance of data to a given task when a model is available, one direct approach is the ”leave-one-out” method [109], where the model’s performance, trained on the entire dataset, is contrasted with the model’s performance when a particular data point is excluded. The difference in performance can then be interpreted as the data point’s relevance for the task. However, computing such values for all data points using this method can be excessively resource-intensive. A more efficient alternative to address this challenge is the use of influence functions [110, 111], which assess the influence of a specific sample on the model’s parameters when that sample’s weight is increased. To further address the complexity of sample interactions, the concept of Data Shapley has emerged [35], drawing inspiration from Shapley values. Data Shapley provides a more comprehensive framework for evaluating the value of individual data points in a dataset and efficient data Shapley approaches have been proposed [112]. However, such methods have not been used for geospatial data so far. In the context of incremental learning in the geospatial domain, Roscher et al. [96] uses the classification model to identify relevant samples based on their ability to help the model to adapt to the new data. More techniques that rely on an existing model are discussed in Section 3.3.
3.2 Data Curation (Step 3)
@@
Accuracy. One of the most commonly used sets of techniques is data-cleaning, which specifically focuses on data accuracy and consistency [18]. Data-only unsupervised techniques are applied before model training and are independent of the learning objective and application. However, as discussed by Ilyas et al. and Neutatz et al. [114, 115], methods that clean the data with respect to the application goals and take the ML model into account show more promising results. Core-set selection techniques such as the one presented by Santos et al. [70] can be considered a data-cleaning approach as the selection process retains the cleanest samples. In their approach, they use clustering for satellite time series to identify samples that are mislabeled or have low accuracy with the goal of removing them from the training set to avoid a decrease in model performance. In some cases, the source of uncertainty is known and can be directly reduced. For instance, data uncertainty may arise from the presence of clouds, and many methods have been developed to automatically remove them [63, 66, 67] (see also techniques that act on diversity and completeness). Northcutt et al. [71], for example, use confident learning [103] to identify and correct label errors. Labels are flagged as noisy when their label uncertainty is high, which means the model is not confident about the label (see Section 4.3). These labeled samples are then iteratively re-labeled using a more accurate labeling function, which can be a human annotator or a more sophisticated model. When multiple labelers provide labels, similar strategies can be used to identify the most likely labels or remove highly confusing samples [26]. Most model-based approaches, however, use data with their quality information during model training instead of cleaning the data and re-training the model in two separate steps (see Model training and data utilization (Step 4)).
@@
Relevance. Curation also considers the relevance of the data. By removing irrelevant data, curation ensures that the data is tailored to specific needs and goals, focusing on the most relevant information. It also naturally reduces the dataset size, improving data efficiency. Core-set selection methods compute the distance between training and test samples and remove those with low proximity. For example, in geospatial air quality estimation, Stadtler et al. [69] demonstrate that removing the 10% of training samples with the lowest proximity to the test samples only slightly decreases test accuracy, as these samples are not relevant for training. However, they assume that the test set is representative enough for the complete data set. A more sophisticated approach is implemented and described in our first experiments (Section 4.2).
@@
3.3 Model training and data utilization (Step 4)
@@
3.4 Model Evaluation (Step 5)
@@ 4 Validation studies
4.1 DFC2020 Land Cover Classification and Experimental Setup
@@
4.2 Study 1: Relevance-based sample weighting to compensate for geographic domain shift
@@
(a) Target accuracy using regular-uniform sampling with relevance-weighted sampling of training data
region B. Anz. Mumbai Mex. C. C. Town B. For. Chab. Kip. R.
uniform 54.73 49.78 60.33 81.11 69.04 25.93 58.63
weighted 55.47 50.78 59.92 80.90 75.35 28.19 59.80
difference 0.74 1.00 -0.41 -0.21 6.30 2.26 1.17
Figure 3: Experiment 1: Sample-wise debiasing geographic domain shift with KLIEP.
@@
Results and Interpretation. We show results in Figs. 3(c), 3(d) and 1. First, Fig. 3(c) shows a world map with average relevance of all images within each source DFC region (scaled circles) with respect to Mumbai. This generally covers Tobler’s [136] intuition that more geographically nearby areas, like Bandar Anzali, are more relevant than distant ones. However, environmental conditions beyond mere geographic distance also contribute to the relevance, as the high importance of Mexico City images shows that it shares a similar arid climate. Second, Fig. 3(d) extends this comparison across all regions and shows the sample-averaged importance scores as edge width and color in a bi-directional graph between all DFC different regions. Here, we can see more generally that nearby (Mumbai $\leftrightarrow$ Bandar Anzali) temperate and forested regions (Black Forest $\leftrightarrow$ Chabaraovsk) and arid regions (Cape Town $\leftrightarrow$ Mexico City $\leftrightarrow$ Mumbai $\leftrightarrow$ Kippa Ring) are more relevant for each other.
Table 1: Target accuracy in [%] using regular-uniform weighting vs. relevance-weighting of training data
@@
Finally, Table 1 quantifies the accuracy benefit of classifying on a sample-weighted training source dataset with respect to each target region in the columns. The row “uniform weighting” shows classification results on the random forest trained on regular uniformly weighted training data, while “relevance weighting” reports the accuracy on KLIEP-weighted training samples. The “difference” rows indicate in which regions the de-biasing was beneficial (positive) or detrimental (negative) to the final target accuracy. Here, the weighting positively affected five of the seven DFC regions up to 6.3%. Two regions were classified at lower accuracy (up to -0.41%). That this procedure can also have detrimental effects in same case illustrates the inherent challenge of deriving the relevance of samples and that negative transfer can occur and deteriorate performance.
@@
4.3 Study 2: Pruning noisy data with confident learning
@@
Table 2: Quantitative results of study 2. Model performance when training on the entire training dataset (baseline) and after pruning noisy training samples (confident learning).
@@
Results and Interpretation. The confident learning method flagged 123 samples in the DFC training set as having potential label issues. Six of these samples are presented in Fig. 4(a), illustrating instances where the model disagreed with the given label. In some cases, as seen in the lower left image, discrepancies arise from the presence of multiple classes within one patch, while in other cases, such as the top left image, the reason is related to the similarity in appearance between classes. Figure 4(b) shows a histogram of the label quality of all training samples belonging to class “Urban/Built-up”, accompanied by three example images representing different label qualities: one with high label quality, one with medium label quality, and one with low label-quality. These examples show that the label quality tends to degrade as the fraction of “Urban/Built-up” pixels within the patch decreases. To improve the quality of the training data, the samples flagged as noisy were removed. Re-training the model on this pruned dataset resulted in an overall accuracy improvement of +3.45 percentage points compared to the baseline (see Table 2).
@@
4.4 Study 3: Evaluation with slice discovery
@@
Figure 5: The confusion matrices (left: validation set, right: test set) of the model described in Section 4.1 for validation (left) and test set (right).
@@
5 Discussion & Conclusion
Shifting research focus from well-explored model-centric approaches towards the broader aspects of the entire machine learning cycle, especially data-centric learning techniques, is necessary to make further progress towards accurate and generalizable models. Here, we find that so far, only limited research is focused on automated data-centric approaches that explicitly improve the quality criteria that we categorized in Section 2 or utilize them for the model-building process. In this work, we provided an extensive literature review in Section 3 and categorized the techniques according to steps in the machine learning cycle and identified the quality criteria they improve or use. This literature study was complemented by concrete implementation examples in Section 4 to stress the applicability of some of the data-centric methods.
@@ 5 Discussion & Conclusion
Experimental Performance. Another important point of discussion is the effectiveness of current data-centric approaches in experiments. While all proposed methods are well-motivated, the actual improvement in accuracy often falls short of expectations. For example, in Study 1 (Section 4.2), weighting dataset samples regarding their relevance only improves two out of five regions. Similarly, in Study 2 (Section 4.3), the quantitative accuracy gain by removing samples with label noise only results in a 3.5% improvement. However, both approaches reveal underlying patterns in the data that can provide valuable insights on their own. Additionally, in real-world applications, accuracy is not the sole factor of interest. Factors such as reliability, precision, robustness, and adaptability are often equally important. Furthermore, a major goal is to avoid the negative effects caused by low-quality data [17]. We further want to point out that the used methods were chosen due to their already proven performance in diverse fields. However, the success of data-centric techniques also depend on the chosen model, therefore we encourage considering the whole machine learning cycle, including model- and data-centric aspects.

### v4

# Better, Not Just More: Data-Centric Machine Learning for Earth Observation

@@ Abstract
Index Terms:
data-centric machine learning, data curation, data utilization, data quality.
I Introduction
@@
Figure 1: Steps of the machine learning cycle. Each step highlights one way to interact with data and its quality, and each step can employ multiple techniques to perform the interaction (see Section III). It involves problem definition, data creation and curation, model training and evaluation, and final deployment, which feeds back into a modified problem definition. Model-centric learning focuses primarily on model training and evaluation, while data-centric learning involves algorithms covering data curation, creation, specific training strategies, evaluation, and deployment feedback. The considered quality criteria are diversity and completeness, accuracy, consistency, unbiasedness, and relevance, see Section II).
@@
The article is organized as follows: It first clarifies the terminology, definitions, and goals of data-centric and model-centric learning (Section II) and then provides a comprehensive discussion of data-centric machine learning and its steps, with a focus on the analysis and interpretation of geospatial data. Additionally, it presents a systematic review of data-centric machine learning techniques tailored explicitly to geospatial data (Section III), facilitating the learning and evaluation of high-performing machine learning models (see Fig. 1). The paper is supported by three experiments using data-centric machine learning techniques, presented in Section IV.
II Definitions and Terminology
@@
II-A Machine learning cycle
@@
II-B High-quality criteria for geospatial data
@@
III Data-centric machine learning techniques
@@
Step 5: evaluationuse quality informationStep 4: model traininguse quality informationStep 3: curationact on qualityStep 2: creationact on qualitydiversity & completenessaccuracyconsistencyunbiasednessrelevance Train/eval splits [40, 41] Resampling techniques [42] Slicing [43] Out-of-distribution evaluation [44, 45] Adversarial attacks [46] Sensitivity analysis [47, 48] Curriculum learning [49, 50, 51, 52, 53, 54] Quality-based weighted loss [55, 56] Pre-training [57, 58, 59, 60, 61] Continual learning [62] Data imputation and inpainting [63, 64, 65, 66, 67, 68] Core-set selection [69, 70] Label noise reduction and confident learning[71] Multi-labeler error detection [72] Bias mitigation [73] Dimensionality reduction [74, 75, 76] Self-training [77] Augmentation [78, 79, 80, 81] Collection and sampling strategies [82, 83, 84, 58, 85, 86, 87, 88, 85, 89, 90] Active learning [91, 92, 93] Confident learning [94] Multi-labeler agreement assessment [33, 95] Bias detection and mitigation [27] Relevance and influence assessment [96, 90]
@@
III-A Data Creation (Step 2)
@@
To quantify the relevance of data to a given task when a model is available, one direct approach is the ”leave-one-out” method [109], where the model’s performance, trained on the entire dataset, is contrasted with the model’s performance when a particular data point is excluded. The difference in performance can then be interpreted as the data point’s relevance for the task. However, computing such values for all data points using this method can be excessively resource-intensive. A more efficient alternative to address this challenge is the use of influence functions [110, 111], which assess the influence of a specific sample on the model’s parameters when that sample’s weight is increased. To further address the complexity of sample interactions, the concept of Data Shapley has emerged [35], drawing inspiration from Shapley values. Data Shapley provides a more comprehensive framework for evaluating the value of individual data points in a dataset and efficient data Shapley approaches have been proposed [112]. However, such methods have not been used for geospatial data so far. In the context of incremental learning in the geospatial domain, Roscher et al. [96] uses the classification model to identify relevant samples based on their ability to help the model to adapt to the new data. More techniques that rely on an existing model are discussed in Section III-C.
III-B Data Curation (Step 3)
@@
Accuracy. One of the most commonly used sets of techniques is data-cleaning, which specifically focuses on data accuracy and consistency [18]. Data-only unsupervised techniques are applied before model training and are independent of the learning objective and application. However, as discussed by Ilyas et al. and Neutatz et al. [114, 115], methods that clean the data with respect to the application goals and take the ML model into account show more promising results. Core-set selection techniques such as the one presented by Santos et al. [70] can be considered a data-cleaning approach as the selection process retains the cleanest samples. In their approach, they use clustering for satellite time series to identify samples that are mislabeled or have low accuracy with the goal of removing them from the training set to avoid a decrease in model performance. In some cases, the source of uncertainty is known and can be directly reduced. For instance, data uncertainty may arise from the presence of clouds, and many methods have been developed to automatically remove them [63, 66, 67] (see also techniques that act on diversity and completeness). Northcutt et al. [71], for example, use confident learning [103] to identify and correct label errors. Labels are flagged as noisy when their label uncertainty is high, which means the model is not confident about the label (see Section IV-C). These labeled samples are then iteratively re-labeled using a more accurate labeling function, which can be a human annotator or a more sophisticated model. When multiple labelers provide labels, similar strategies can be used to identify the most likely labels or remove highly confusing samples [26]. Most model-based approaches, however, use data with their quality information during model training instead of cleaning the data and re-training the model in two separate steps (see Model training and data utilization (Step 4)).
@@
Relevance. Curation also considers the relevance of the data. By removing irrelevant data, curation ensures that the data is tailored to specific needs and goals, focusing on the most relevant information. It also naturally reduces the dataset size, improving data efficiency. Core-set selection methods compute the distance between training and test samples and remove those with low proximity. For example, in geospatial air quality estimation, Stadtler et al. [69] demonstrate that removing the 10% of training samples with the lowest proximity to the test samples only slightly decreases test accuracy, as these samples are not relevant for training. However, they assume that the test set is representative enough for the complete data set. A more sophisticated approach is implemented and described in our first experiments (Section IV-B).
@@
III-C Model training and data utilization (Step 4)
@@
III-D Model Evaluation (Step 5)
@@ IV Validation studies
IV-A DFC2020 Land Cover Classification and Experimental Setup
@@
IV-B Study 1: Relevance-based sample weighting to compensate for geographic domain shift
@@
Figure 3: Study 1. Analyzing geographic domain shift with KLIEP to derive a relevance-weighting scheme for model training.
@@
Results and Interpretation. We show results in Figs. 3(c), 3(d) and I. First, Fig. 3(c) shows a world map with average relevance of all images within each source DFC region (scaled circles) with respect to Mumbai. This generally covers Tobler’s [136] intuition that more geographically nearby areas, like Bandar Anzali, are more relevant than distant ones. However, environmental conditions beyond mere geographic distance also contribute to the relevance, as the high importance of Mexico City images shows that it shares a similar arid climate. Second, Fig. 3(d) extends this comparison across all regions and shows the sample-averaged importance scores as edge width and color in a bi-directional graph between all DFC different regions. Here, we can see more generally that nearby (Mumbai $\leftrightarrow$ Bandar Anzali) temperate and forested regions (Black Forest $\leftrightarrow$ Chabaraovsk) and arid regions (Cape Town $\leftrightarrow$ Mexico City $\leftrightarrow$ Mumbai $\leftrightarrow$ Kippa Ring) are more relevant for each other.
Table I: Target accuracy in [%] using regular-uniform weighting vs. relevance-weighting of training data
@@
Finally, Table I quantifies the accuracy benefit of classifying on a sample-weighted training source dataset with respect to each target region in the columns. The row “uniform weighting” shows classification results on the random forest trained on regular uniformly weighted training data, while “relevance weighting” reports the accuracy on KLIEP-weighted training samples. The “difference” rows indicate in which regions the de-biasing was beneficial (positive) or detrimental (negative) to the final target accuracy. Here, the weighting positively affected five of the seven DFC regions up to 6.3%. Two regions were classified at lower accuracy (up to -0.41%). That this procedure can also have detrimental effects in same case illustrates the inherent challenge of deriving the relevance of samples and that negative transfer can occur and deteriorate performance.
@@
IV-C Study 2: Pruning noisy data with confident learning
@@
Table II: Quantitative results of study 2. Model performance when training on the entire training dataset (baseline) and after pruning noisy training samples (confident learning).
@@
Results and Interpretation. The confident learning method flagged 123 samples in the DFC training set as having potential label issues. Six of these samples are presented in Fig. 4(a), illustrating instances where the model disagreed with the given label. In some cases, as seen in the lower left image, discrepancies arise from the presence of multiple classes within one patch, while in other cases, such as the top left image, the reason is related to the similarity in appearance between classes. Figure 4(b) shows a histogram of the label quality of all training samples belonging to class “Urban/Built-up”, accompanied by three example images representing different label qualities: one with high label quality, one with medium label quality, and one with low label-quality. These examples show that the label quality tends to degrade as the fraction of “Urban/Built-up” pixels within the patch decreases. To improve the quality of the training data, the samples flagged as noisy were removed. Re-training the model on this pruned dataset resulted in an overall accuracy improvement of +3.45 percentage points compared to the baseline (see Table II).
@@
IV-D Study 3: Evaluation with slice discovery
@@
Figure 5: The confusion matrices (left: validation set, right: test set) of the model described in Section IV-A for validation (left) and test set (right).
@@
V Discussion & Conclusion
Shifting research focus from well-explored model-centric approaches towards the broader aspects of the entire machine learning cycle, especially data-centric learning techniques, is necessary to make further progress towards accurate and generalizable models. Here, we find that so far, only limited research is focused on automated data-centric approaches that explicitly improve the quality criteria that we categorized in Section II or utilize them for the model-building process. In this work, we provided an extensive literature review in Section III and categorized the techniques according to steps in the machine learning cycle and identified the quality criteria they improve or use. This literature study was complemented by concrete implementation examples in Section IV to stress the applicability of some of the data-centric methods.
@@ V Discussion & Conclusion
Experimental Performance. Another important point of discussion is the effectiveness of current data-centric approaches in experiments. While all proposed methods are well-motivated, the actual improvement in accuracy often falls short of expectations. For example, in Study 1 (Section IV-B), weighting dataset samples regarding their relevance only improves two out of five regions. Similarly, in Study 2 (Section IV-C), the quantitative accuracy gain by removing samples with label noise only results in a 3.5% improvement. However, both approaches reveal underlying patterns in the data that can provide valuable insights on their own. Additionally, in real-world applications, accuracy is not the sole factor of interest. Factors such as reliability, precision, robustness, and adaptability are often equally important. Furthermore, a major goal is to avoid the negative effects caused by low-quality data [17]. We further want to point out that the used methods were chosen due to their already proven performance in diverse fields. However, the success of data-centric techniques also depend on the chosen model, therefore we encourage considering the whole machine learning cycle, including model- and data-centric aspects.